# Device-Independent Private Quantum Randomness Beacon

### v1

@@ III Single-round analysis
The guessing probability can then be bounded using the Navascués-Pironio-Acín (NPA) hierarchy [npa]. We shall solve the following optimisation problem
@@
We note that it is also straightforward to modify the optimisation problems for protocols that monitor linear combinations of $\{p(ab|xy,s=0)\}_{abxy}$ and linear combinations of $\{p(ac|xz,s=1)\}_{acxz}$. From the guessing probability, we can obtain an affine min-tradeoff function by simply linearising the logarithmic function by taking a tangent at any given point in the interval $[0,1]$. We also note that it is possible to tighten our bound by using the technique presented in Refs. [tan2024entropy, deloison2025device] which is based on the Brown-Fawzi-Fawzi bound [brown2024device], however, this will require us to solve a bigger SDP.

### v2

@@ V.3 Single-round analysis
The guessing probability can then be bounded using the Navascués-Pironio-Acín (NPA) hierarchy [32]. We shall solve the following optimisation problem
@@
We note that it is also straightforward to modify the optimisation problems for protocols that monitor linear combinations of $\{p(ab|xy,s=0)\}_{abxy}$ and linear combinations of $\{p(ac|xz,s=1)\}_{acxz}$. From the guessing probability, we can obtain an affine min-tradeoff function by simply linearising the logarithmic function by taking a tangent at any given point in the interval $[0,1]$. We also note that it is possible to tighten our bound by using the technique presented in Refs. [33, 34] which is based on the Brown-Fawzi-Fawzi bound [35], however, this will require us to solve a bigger SDP.